# Prime Integer Matrices

### v1

@@ Abstract
This paper introduces prime integer matrices and its properties. It provides a simple way to construct families of pairwise co-prime integer matrices, that may have applications in multidimensional co-prime sensing and multidimensional Chinese remainder theorem.
@@ I Introduction
The CRT has been recently generalized to general integer matrices of any form and any dimension in [14, 15] to reconstruct integer vectors from their vector remainders modulo several integer matrices called matrix moduli, i.e., multidimensional CRT (MD-CRT). Similar to the conventional CRT for integers, to have a large range of uniquely reconstructable integer vectors from their vector remainders, it is critical to have a large size set of pairwise co-prime integer matrices as matrix moduli. Also, co-prime integer matrices have applications in multidimensional sparse sensing [8]. Unfortunately, due to the non-commutativity of matrices, it is not obvious to construct a family of pairwise co-prime integer matrices of any dimension, although co-primality of integer matrices has been studied in [7] mostly for dimensions $2$ and $3$. Interestingly, a family of pairwise co-prime integer matrices was most recently obtained in [16] for any dimension.
@@
It is known that any set of prime integers are pairwise co-prime. This holds for any prime algebraic integers as well. Therefore, it is usually an easy way to construct families of pairwise co-prime elements by simply constructing families of prime elements, such as prime integers. This motivates us to study prime integer matrices, despite its other potential applications as prime elements.
@@
In this paper, we define prime integer matrices and study their properties. We show that an integer matrix (square matrix) is prime if and only if its determinant absolute value is a conventional prime integer. We also show that, similar to prime integers, two different prime integer matrices are co-prime. We then present the Hermite normal forms [13] for prime integer matrices, and show that two prime integer matrices with the same determinant absolute value but different Hermite normal forms are co-prime. Thus, it is easy to construct families of pairwise co-prime integer matrices by simply contructing families of prime integer matrices with different determinant absolute values that are prime and/or with the same determinant absolute value that is prime but different Hermite normal forms. It is interesting to note that none of the pairwise co-prime integer matrices in the construction in [16] is prime when the dimension is above $1$. Also note that a different kind of primality for matrices was studied in [11, 12] for a given set of matrices.
@@
Since Gaussian integers are special $2\times 2$ integer matrices, we find that Gaussian primes and prime $2\times 2$ integer matrices are not exactly the same. When both real and imaginary parts of Gaussian integers are not 00, Gaussian primes and prime $2\times 2$ integer matrices are the same. Otherwise, Gaussian primes may not be prime $2\times 2$ integer matrices. It is well-known that any Gaussian integer has a unique Gaussian prime factorization [2]. We show that an integer matrix can be unqiuely factorized to a product of prime integer matrices when the order of the prime integer matrices in the product is not considered. Note that the order of primes in a product is not a problem if the primes are communtative, such as Gaussian primes (or their equivalent $2\times 2$ integer matrices).
@@
This paper is organized as follows. In Section II, we introduce prime integer matrices and present their properties and Hermite normal forms. In Section III, we present some connections between Gaussian primes and prime $2\times 2$ integer matrices. In Section IV, we conclude this paper.
@@ II Definitions and Properties of Prime Integer Matrcies
We next introduce prime matrices.
@@
A nonsingular matrix $A$ is called left prime (or simply prime) if it cannot be factorized to a product of two non-unimodular matrices. If $A$ is prime, then $A$ is called a prime matrix.
@@
A matrix is prime if and only if its determinant absolute value is a conventional prime integer.
Proof. If matrix $A$ is not prime, then there exist two non-unimodular matrices $B$ and $C$ such that $A=BC$. Since $|\mbox{det}(A)|=|\mbox{det}(B)|\cdot|\mbox{det}(C)|$, the determinant absolute value of matrix $A$ is not prime. This proves the sufficiency.
We next prove the necessity. If the determinant absolute value of matrix $A$ is not a prime integer, i.e., $|\mbox{det}(A)|=ab$ for two positive integers $a,b>1$. Let the Smith form of matrix $A$ be $A=U\Lambda V$ where matrices $U$ and $V$ are unimodular and matrix $\Lambda$ is a diagonal integer matrix. Thus, $|\mbox{det}(A)|=|\mbox{det}(\Lambda)|$. Since $|\mbox{det}(A)|=ab$, diagonal matrix $\Lambda$ can be written as a product of two diagonal integer matrices as $\Lambda=\Lambda_{1}\Lambda_{2}$ for diagonal matrices $\Lambda_{1}$ with $|\mbox{det}(\Lambda_{1})|=a$ and $\Lambda_{2}$ with $|\mbox{det}(\Lambda_{2})|=b$. Thus, we have $A=U\Lambda_{1}\Lambda_{2}V$ and matrices $U\Lambda_{1}$ and $\Lambda_{2}V$ are not unimodular. This proves that $A$ is not prime and thus proves the necessity. q.e.d.
@@
We next provide a complete answer to this question by using Hermite normal form of integer matrices (Chapter 14 of [13]) that says that any matrix has a unique Hermite normal form. Combined with the primality of a matrix with the result in Corollary 1, we have the following proposition.
@@
All prime matrices that have the same determinant absolute value but different Hermite normal forms are different and pairwise co-prime.
@@
As we can see from the Hermite normal form $H$ in (2), for any prime integer $\lambda_{D}>1$, its different positions $m_{0}$ on the diagonal and different values of $a_{1},...,a_{k}$ all produce different Hermite normal forms and therefore different prime matrices, i.e., they are all pairwise co-prime, from Props. 1 and 4. Thus, from Corollary 2 and Prop. 4, one is able to construct many families of pairwise co-prime matrices with different determinant absolute values that are prime and with different Hermite normal forms but the same determinant absolute value that is a prime. Interestingly, a family of pairwise co-prime integer matrices of any dimension, none of which is prime when the dimension is more than $1$, was obtained in [16].
@@ III Connection with Gaussian Integers and Gaussian Primes
A matrix $A$ can be uniquely factorized to a product of prime matrices called prime factors, if the order the prime factors in the product is not considered.
@@
Although the above connection is only studied for Gaussian integers, similar studies may apply to other algebraic integers in other algebraic number fields.
@@ IV Conclusion
In this paper, we have introduced and characterized prime integer matrices and their Hermite normal forms. An integer matrix is prime if and only if its determinant absolute value is a conventional prime integer. Two prime integer matrices with different determinant absolute values that are prime are co-prime. Two prime integer matrices with the same determinant absolute value that is prime but different Hermite normal forms are co-prime. From these results, we are able to easily construct pairwise co-prime integer matrix families that have applications in multidimensional CRT and multidimensional sparse sensing. We have also obtained the connections between prime integer matrices and Gaussian primes. While the unique factorization of a Gaussian integer as a product of Gaussian primes holds, we have shown that any integer matrix can be uniquely factorized to a product of prime matrices if the order of the prime matrices in the product is not considered.

### v2

# Prime and Co-prime Integer Matrices

@@ Abstract
This paper investigates prime and co-prime integer matrices and their properties. It characterizes all pairwise co-prime integer matrices that are also prime integer matrices. This provides a simple way to construct families of pairwise co-prime integer matrices, that may have applications in multidimensional co-prime sensing and multidimensional Chinese remainder theorem.
@@ I Introduction
Prime integers have also been generalized to prime integer matrices in [2] where it obtains that an integer matrix is prime if and only if its determinant absolute value is a conventional prime integer. Recently, CRT has been generalized to general integer matrices of any form and any dimension in [15, 16] to reconstruct integer vectors from their vector remainders modulo several integer matrices called matrix moduli, i.e., multidimensional CRT (MD-CRT). Similar to the conventional CRT for integers, to have a large range of uniquely reconstructable integer vectors from their vector remainders, it is critical to have a large size set of pairwise co-prime integer matrices as matrix moduli. Also, co-prime integer matrices have applications in multidimensional sparse sensing [9]. Unfortunately, due to the non-commutativity of matrices, it is not obvious to construct a family of pairwise co-prime integer matrices of any dimension, although co-primality of integer matrices has been studied in [8] mostly for dimensions $2$ and $3$. Interestingly, a family of pairwise co-prime integer matrices was most recently obtained in [17] for any dimension.
@@
It is known that any set of prime integers are pairwise co-prime. This holds for any prime algebraic integers as well. Therefore, it is usually an easy way to construct families of pairwise co-prime elements by simply constructing families of prime elements, such as prime integers. This motivates us to use prime integer matrices to characterize and construct pairwise co-prime integer matrices that are also prime.
@@
In this paper, we first recall prime integer matrices and some of their basic properties presented in [2]. We then investigate co-prime and pairwise co-prime integer matrices. We show that, similar to prime integers, two different prime integer matrices are co-prime. We then present the Hermite normal forms [14] for prime integer matrices, and show that two prime integer matrices with the same determinant absolute value but different Hermite normal forms are co-prime. Thus, it is easy to construct families of pairwise co-prime integer matrices by simply contructing families of prime integer matrices with different determinant absolute values that are prime and/or with the same determinant absolute value that is prime but different Hermite normal forms. This characterizes all pairwise co-prime integer matrices that are all prime. It is interesting to note that none of the pairwise co-prime integer matrices in the construction in [17] is prime when the dimension is above $1$. Also note that a different kind of primality for matrices was studied in [12, 13] for a given set of matrices.
@@
Since Gaussian integers are special $2\times 2$ integer matrices, interestingly, Gaussian primes and prime $2\times 2$ integer matrices are not exactly the same. When both real and imaginary parts of Gaussian integers are not 00, Gaussian primes and prime $2\times 2$ integer matrices are the same. Otherwise, Gaussian primes may not be prime $2\times 2$ integer matrices. However, two Gaussian integers are co-prime if and only if their equivalent $2\times 2$ integer matrices are co-prime. It is well-known that any Gaussian integer has a unique Gaussian prime factorization [3]. In contrast, an integer matrix can be unqiuely factorized to a product of prime integer matrices when the order of the prime integer matrices in the product is not considered. Note that the order of primes in a product is not a problem if the primes are communtative, such as Gaussian primes (or their equivalent $2\times 2$ integer matrices).
@@
This paper is organized as follows. In Section II, we study prime integer matrices and their properties. We show their Hermite normal forms and characterize all pairwise co-prime integer matrices that are all prime. In Section III, we present some connections between Gaussian primes and prime $2\times 2$ integer matrices. In Section IV, we conclude this paper.
@@ II Definitions and Properties of Prime Integer Matrcies
We next introduce prime matrices [2], Chapter 14.
@@
[2] A nonsingular matrix $A$ is called left prime (or simply prime) if it cannot be factorized to a product of two non-unimodular matrices. If $A$ is prime, then $A$ is called a prime matrix.
@@
[2] A matrix is prime if and only if its determinant absolute value is a conventional prime integer.
Its proof is not hard and can be found in [2].
@@
We next provide a complete answer to this question by using Hermite normal form of integer matrices (Chapter 14 of [2] and Chapter 14 of [14]) where is says that any matrix has a unique Hermite normal form. Combined with the primality of a matrix with the result in Corollary 1, we have the following proposition.
@@
All prime matrices that have the same determinant absolute value but different Hermite normal forms (2) are different and pairwise co-prime.
@@
As we can see from the Hermite normal form $H$ in (2), for any prime integer $\lambda_{D}>1$, its different positions $m_{0}$ on the diagonal and different values of $a_{1},...,a_{k}$ all produce different Hermite normal forms and therefore different prime matrices, i.e., they are all pairwise co-prime, from Props. 1 and 4. Thus, from Corollary 2 and Prop. 4, one is able to construct many families of pairwise co-prime matrices with different determinant absolute values that are prime and with different Hermite normal forms but the same determinant absolute value that is a prime. The above results have also characterized all pairwise co-prime integer matrices that are all prime. Interestingly, a family of pairwise co-prime integer matrices of any dimension, none of which is prime when the dimension is more than $1$, was obtained in [17].
@@ III Connection with Gaussian Integers and Gaussian Primes
A non-singular matrix $A$ can be uniquely factorized to a product of prime matrices called prime factors, if the order the prime factors in the product is not considered.
@@
A similar integer matrix factorization can be found in [2]. Although the above connection is only studied for Gaussian integers, similar studies may apply to other algebraic integers in other algebraic number fields.
@@ IV Conclusion
In this paper, we have applied prime integer matrices to construct and characterize all pairwise co-prime integer matrices that are all prime. For instance, two prime integer matrices with different determinant absolute values that are prime are co-prime, and two prime integer matrices with the same determinant absolute value that is prime but different Hermite normal forms are co-prime. These pairwise co-prime integer matrices may have applications in multidimensional CRT and multidimensional sparse sensing. We have also investigated the connections between prime integer matrices and Gaussian primes.